Operations Report – March 9th

Crew 206 Operations Report 09 – MAR – 2019

SOL: 13

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 13,5 hours

Solar— SOC % – Turned off at 80%, Turned on at 73% (18h44)

Diesel Reading – 60%
Propane Reading – 76%

Ethanol Free Gasoline – 2 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 375 gallons

Water differential (static tank) – 13 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 226 gallons

Static to Loft Pump used – Yes

Water Meter: 0141711.2 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional
Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system tested.
Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Used
Hours: 78.2 hours

Beginning charge: 90%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Used
Hours: 72,2 hours

Beginning charge: 100%

Ending charge: 72 %

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is empty, it looks like the pipeline has a leak. Atila is working on it; more information will follow.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used by Staff.

Staff Crew Car used and why, where? Off-Campus

General notes and comments: Nothing to report.

Summary of internet: Internet access fixed! Thank you again, Bernard.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing more to report.